# ZMM-TTS: Zero-shot Multilingual and Multispeaker Speech Synthesis Conditioned on Self-supervised Discrete Speech Representations

### v1

@@ IV-D1 Subjective Evaluation Methods
We synthesized sentences from test sets in six languages for each TTS system. For each language, we select 6 (3 female, 3 male) seen speakers and 6 (3 female, 3 male) unseen speakers for the test set, and two sentences for each speaker. 10 native listeners per language participated in our listening tests. Listeners evaluated one test utterance at a time, initially assigning a rating on a Likert scale from 1 to 5 for the Mean Opinion Score (MOS) to assess naturalness. Subsequently, they provided ratings for speaker similarity concerning a reference utterance using a Differential MOS (DMOS) scale, ranging from 1 (indicating a different speaker) to 5 (indicating the same speaker). Reference utterances were chosen randomly from the original speech of the target speaker. Due to the high cost of subjective evaluation, only the original YourTTS is used as the YourTTS related baseline to test MOS and DMOS. Our preliminary experimental results also demonstrated that YourTTS achieves better performance than YourTTSE and YourTTSW.
@@ IV-D2 Objective Evaluation Methods
We also objectively evaluate performance by measuring the intelligibility of speech content using an ASR algorithm. We synthesized 2,000 sentences (1,000 with seen and unseen speaker embeddings, respectively) for each language. It is important to note that the language of the target speaker and text are always consistent. Although our model also has the ability of cross-lingual synthesis, considering that there is no clear standard for the evaluation of cross-language synthesis, this paper only focuses on the intra-lingual scenarios.
@@ V Results and analysis
The evaluation results for speech naturalness and speaker similarity are shown in Tables IV and V for seen and unseen target speakers, respectively.
@@ V-A1 Comparison between FSM and ZMM-TTS models
As a currently popular method based on FastSpeech and HiFi-GAN, the baseline FastSpeech-based FSM model has achieved SOTA on many monolingual synthesis tasks, while obtaining the worst results in speech naturalness in our multilingual datasets. This result implies that for FSM models based on Mel spectrograms, the sound quality of synthesized audio may be limited due to the sound quality of the training data itself. On the other hand, since the intermediate features used by our ZMM-TTS model are extracted from SSL models trained on large-scale non-ideal data, the proposed system is considered to be less sensitive to the sound quality of the training data. The reason why the proposed method ZMTTS-2x has higher MOS values than ground-truth audio on seen (Spanish) and unseen speakers (French and Swedish) would also be due to the fact that the sound quality of the ground-truth audio is not high.
In terms of speaker similarity, our proposed ZMM-TTS system achieves better DMOS than the FSM-based system. The speaker encoder used in ZMM-TTS and FSM-based systems is the same. Also, the difference between the DMOS of ZMM-TTS and FSM-based systems is smaller than their difference in MOS. This suggests that the improvement in naturalness using features extracted from SSL models is more prominent than the improvement in speaker similarity because discrete features may contain less speaker information.
@@ V-A2 Effectiveness of VQ in vec2wav
Another finding is that adding the vector quantization module in the process of converting the Mel spectrograms or discrete representations to a waveform can improve the naturalness of the audio. Results in Tables IV and V show that FSM2 outperformed FSM1 on MOS. One potential reason is that the discrete representations in vec2wavVQ (Mel) may reduce some redundant information, such as noise, to some extent. Similar findings are observed when comparing ZMM-TTS1 and ZMM-TTS2. ZMM-TTS2 has an additional VQ module and converts discrete representations directly to speech without an independent vocoder. Directly mapping discrete representations to a waveform instead of a Mel spectrogram prevents error propagation caused by inaccurate Mel spectrogram prediction in the vec2mel stage and hence, like the result of FSM1 and FSM2, multi-stage discrete representations learned by additional VQ may also be helpful for improving sound quality.
In terms of speaker similarity, the use of VQ in vec2wav does not always result in an improvement. We also find that MOS is not always correlated with DMOS. For example, as shown in the French results in Table IV, the naturalness of ZMM-TTS2c is significantly higher than that of ZMM-TTS1c (4.90 vs 2.90), while their DMOS values are very similar (4.27 vs 4.28).
@@ V-A3 Comparison between YourTTS and ZMM-TTS models
On speech naturalness, the YourTTS method outperforms the two-stage FSM baseline while remaining worse than two of our proposed systems, ZM-TTS2c and ZM-TTS2x. We observed that the YourTTS model exhibits instability in its stochastic duration predictor, resulting in the production of unnatural durations for certain speakers and sentences.
On speaker similarity, ZM-TTS2c and ZM-TTS2x has better performance than YourTTS systems for the seen speaker condition in five languages (en, fr, ge, pt, and sw). However, under unseen conditions, YourTTS is better than our method in terms of DMOS in English, French, German and Spanish. This will be analyzed in detail in the next subsection.
@@ V-A4 Comparison between different input representations and languages
Interestingly, determining which is better for speech naturalness, character-based inputs or the pre-trained phoneme representation based inputs, depends on the specific language and synthesis systems. FSM1x and FSM2x always have a higher MOS than FSM1c and FSM2c among the FSM-based systems. Utilizing pre-trained phoneme representations as input and XPhoneBERT as the text encoder can enhance the quality of synthesized speech. However, for our proposed ZMM-TTS system, better naturalness is not always achieved with ZMM-TTS1x and ZMM-TTS2x. For example, in Table V, comparing the MOS of ZMM-TTS2x and ZMM-TTS2c, ZMM-TTS2x has achieved better results in English, French, and Portuguese, while ZMM-TTS2c has achieved better results in other languages. Furthermore, we observed that the MOS of ZMM-TTS2x is significantly better than ZMM-TTS2c in English. This may be because XphoneBERT was pre-trained on a dataset with more English text than other languages. We will investigate how the amount of pre-training text affects XphoneBERT’s representation performance in the future.
@@ V-A5 Comparison between seen and unseen speakers
An interesting result is that the naturalness of the synthesized speech of unseen speakers is not worse than that of seen speakers. However, there was still a gap between seen and unseen speakers in terms of DMOS for most languages. Although the pre-training process for the speaker encoder involves more than 7,000 speakers, the TTS system is only trained on a few hundred speakers. It is still challenging to achieve the ideal zero-shot performance under training with limited multilingual data.
@@ V-B Objective Evaluation Results
TABLE VI: CER (%) of ground-truth recordings (GT) and synthesized audio samples from various models.
@@ V-B1 SECS analysis
Firstly, we found very strong correlations ($r=0.807$), using Pearson correlation coefficient (PCC), between the SECS and the DMOS values. Compared to the FSM-based model, ZMM-TTS performs better in both SECS and DMOS metrics for both seen and unseen speakers. Compared to the YourTTS model, ZMM-TTS achieves similar SECS for seen speakers, but there are differences for unseen speakers, particularly English speakers. One possible reason for this difference is that the original implementation of YourTTS utilizes SCL to enable generalization in characteristics of unseen speakers. The different SECS results of YourTTSE and YourTTSW also demonstrate the impact of SCL.
To investigate whether the difference is related to language and the impact of two speaker encoders, we plotted speaker embeddings of seen and unseen speakers from natural speech via T-SNE[65]. Fig. 5 presents the T-SNE visualization and shows the following: Several unseen speakers are not close to the seen speakers in the training set. For example, there are two obvious outliers in the ECAPA-TDNN speaker embeddings of the English unseen speakers. We conducted a study to calculate the SECS values of the 6 unseen English speakers whose speech was synthesized by the ZMM-TTS1c system. We found two English outlier points in Figure 5(a) with SECS values of 0.766 and 0.644, which were significantly lower than the SECS value for the speaker who was closer to the seen English speakers. This explains why the speaker similarity of our proposed ZMM-TTS1c for English unseen speakers is significantly lower than that of YourTTS.
Although there are many different types of SOTA neural speaker embeddings, currently, ZMM-TTS has only attempted using a popular speaker representation from ECAPA-TDNN, and we will explore the performance with different speaker embeddings such as H/ASP and speaker consistency loss in future work.
@@ V-B2 CER analysis
Table VI summarizes the obtained CER. The first finding is that the FSM1 and FSM2 baseline models perform better, although their sound quality is worse than the other systems. This demonstrates that the speech generated using Mel spectrograms as intermediate features is intelligible, but not as natural sounding as when using SSL representations. Additionally, the speech synthesized through the Mel spectrograms tends to over-smooth the high-frequency region more than through SSL representations.
We also find that the CER is not always correlated with the MOS values. Although YourTTS CER is worse, the sound quality is obviously better compared to the FSM-based model. Furthermore, YourTTS CER in Portuguese and Swedish is significantly worse than in other languages. This also explains why the YourTTS system is significantly worse than our ZMM-TTS on MOS in these two languages. YourTTS only uses the character transcriptions rather than phonemes, which makes it more prone to mispronunciation issues as mentioned in [8].
Besides, we found that the pre-trained phoneme representation input system performance is the best, and IPA is only better than characters in English. A notable result is that compared with FSM1, the CER performance of FSM2 has dropped significantly. For example, sometimes the pronunciation of the word “advice” in a sentence becomes incorrect and sounds more like the word “addressed,” resulting in ASR recognition errors. This shows that learning a discrete representation from Mel spectrograms to speech conversion in vec2wavVQ (Mel) can improve the quality of speech naturalness, but some fine-grained information related to linguistic content may be lost.
@@ VI Low-resource scenarios
In addition to the high-resource languages, we also evaluate the performance of the proposed method in unseen low-resource language TTS scenarios to investigate the impact of the training data size.
@@ VI-A2 Implementation Details
Considering the performance on six high-resource languages, we choose three models FSM2, ZMM-TTS2 and YourTTS to implement on new languages. We have tried using both characters and pre-trained phonemes as input representations.
For the few-shot scenario, we created training sets of different sizes (2.5 minutes of audio, 5 minutes of audio, and 15 minutes of audio). We use limited training data to finetune the model trained on six languages in Section IV. For language embeddings, we reserve a free ID for new languages, which will be trained from scratch using fine-tuning data only. Similarly, since different languages may use different characters, symbol embeddings for any characters unseen during pre-training must also be learned from fine-tuning data only. Specifically, in our experiments, we found that Italian and Polish had 3 and 14 unseen symbols, respectively, compared to the pre-training six languages. All few-shot models are fine-tuned for 10,000 steps with a batch size of 16. In order to compare the performance differences of the same language with high and low resources, we also perform fine-tuning with much larger datasets of 35h for Italian and 21h for Polish.
For the zero-shot scenario, to address the mismatch between language embeddings of seen and unseen languages, we removed the language embedding layer. Therefore, the models tested on the zero-shot scenario are trained as in Section-IV on 6 seen languages but without a language embedding layer. Additionally, we used pre-trained phoneme representations as input, making it easily applicable to many unseen languages.
@@ VI-A3 Evaluation Methods
We kept 100 sentences for the speech naturalness and speaker similarity tests and 1,000 sentences for ASR tests like those mentioned in Section IV-C. Furthermore, we adopted a publicly available automatic MOS (AMOS) prediction model [67] to assess naturalness as in [5].
TABLE VII: Low-resource results. Total training data is 35 and 21 hours for Italian and Polish, respectively.
Italian Polish
Size Method AMOS SECS WER CER AMOS SECS WER CER
Few-shot scenarios
015m FSM2x 2.58 0.85 16.28 3.27 2.11 0.89 18.95 5.24
ZMM-TTS2x 2.97 0.91 18.51 3.92 2.77 0.93 26.53 8.27
FSM2c 2.10 0.78 30.26 7.13 1.45 0.83 65.69 28.36
ZMM-TTS2c 2.86 0.90 26.65 6.07 2.55 0.92 40.49 13.34
005m FSM2x 2.60 0.86 16.49 3.39 2.34 0.90 21.01 6.12
ZMM-TTS2x 2.92 0.90 19.29 4.02 2.69 0.92 29.67 10.01
2.5m FSM2x 2.48 0.84 16.04 3.12 2.22 0.89 31.82 11.29
ZMM-TTS2x 2.76 0.89 20.12 4.38 2.61 0.91 38.60 14.64
Zero-shot scenarios
$0$ FSM2x 2.27 0.73 18.23 4.10 2.33 0.77 41.38 15.93
ZMM-TTS2x 3.27 0.83 21.69 5.11 2.42 0.85 38.89 15.40
High-resource baseline
0001h YourTTS 1.64 0.74 99.21 89.77 1.40 0.65 98.57 88.86
35/21h YourTTS 2.91 0.96 27.18 6.01 1.70 0.71 100.37 89.68
35/21h FSM2x 2.27 0.91 13.53 3.29 2.00 0.88 11.40 2.89
35/21h ZMM-TTS2x 3.02 0.96 15.69 3.76 2.87 0.95 15.71 4.49
35/21h FSM2c 2.17 0.88 15.36 2.47 1.79 0.85 14.98 3.71
35/21h ZMM-TTS2c 2.92 0.96 13.87 2.22 2.70 0.95 11.89 2.42
- GT 2.94 0.99 12.62 3.17 2.82 0.99 10.98 3.95

### v2

@@ IV-D1 Subjective Evaluation Methods
We synthesized sentences from test sets in six languages for each TTS system. For each language, we select 6 (3 female, 3 male) seen speakers and 6 (3 female, 3 male) unseen speakers for the test set, and two sentences for each speaker. Ten native listeners per language participated in our listening tests. Listeners evaluated one test utterance at a time, initially assigning a rating on a Likert scale from 1 to 5 for the Mean Opinion Score (MOS) to assess naturalness. Subsequently, they provided ratings for speaker similarity concerning a reference utterance using a Differential MOS (DMOS) scale, ranging from 1 (indicating a different speaker) to 5 (indicating the same speaker). Reference utterances were chosen randomly from the original speech of the target speaker. Due to the high cost of subjective evaluation, only the original YourTTS is used as the YourTTS related baseline to test MOS and DMOS. Our preliminary experimental results also demonstrated that YourTTS achieves better performance than YourTTSE and YourTTSW.
@@ IV-D2 Objective Evaluation Methods
We also objectively evaluate performance by measuring the intelligibility of speech content using an ASR algorithm. We synthesized 2,000 sentences (1,000 with seen and unseen speaker embeddings, respectively) for each language. Note that the language of the target speaker and text are always consistent. Although our model also has the ability of cross-lingual synthesis, considering that there is no clear standard for the evaluation of cross-language synthesis, this paper only focuses on the intra-lingual scenarios.
@@ V Results and analysis
The evaluation results for speech naturalness and speaker similarity are shown in Tables V and VI for seen and unseen target speakers, respectively.
@@ V-A1 Comparison between FSM and ZMM-TTS models
As a currently popular method based on FastSpeech and HiFi-GAN, the baseline FastSpeech-based FSM model has achieved SOTA on many monolingual synthesis tasks while obtaining the worst results in speech naturalness in our multilingual datasets. This result indicates that for FSM models based on Mel spectrograms, the sound quality of synthesized audio may be limited due to the sound quality of the training data itself. On the other hand, since the intermediate features used by our ZMM-TTS model are extracted from SSL models trained on large-scale non-ideal data, the proposed system is considered to be less sensitive to the sound quality of the training data. The reason that the proposed method ZMTTS-2x has higher MOS values than ground-truth audio on seen (Spanish) and unseen speakers (French and Swedish) would also be due to the fact that the sound quality of the ground-truth audio is not high.
In terms of speaker similarity, our proposed ZMM-TTS system achieves better DMOS than the FSM-based system. The speaker encoder used in the ZMM-TTS and FSM-based systems is the same. Also, the difference between the DMOS of the ZMM-TTS and FSM-based systems is smaller than their difference in MOS. This suggests that the improvement in naturalness using features extracted from SSL models is more prominent than the improvement in speaker similarity because discrete features may contain less speaker information.
@@ V-A2 Effectiveness of VQ in vec2wav
Another finding is that adding the VQ module in the process of converting the Mel spectrograms or discrete representations to a waveform can improve the naturalness of the audio. The results in Tables V and VI show that FSM2 outperformed FSM1 on MOS. One potential reason is that the discrete representations in vec2wavVQ (Mel) may reduce some redundant information, such as noise, to some extent. Similar findings are observed when comparing ZMM-TTS1 and ZMM-TTS2. ZMM-TTS2 has an additional VQ module and converts discrete representations directly to speech without an independent vocoder. Directly mapping discrete representations to a waveform instead of a Mel spectrogram prevents error propagation caused by inaccurate Mel spectrogram prediction in the vec2mel stage and hence, like the result of FSM1 and FSM2, multi-stage discrete representations learned by additional VQ may also be helpful for improving sound quality.
In terms of speaker similarity, the use of VQ in vec2wav does not always result in an improvement. We also find that MOS is not always correlated with DMOS. For example, as shown in the French results in Table V, the naturalness of ZMM-TTS2c is significantly higher than that of ZMM-TTS1c (4.20 vs 2.90), while their DMOS values are very similar (4.27 vs 4.28).
@@ V-A3 Comparison between YourTTS and ZMM-TTS models
On speech naturalness, the YourTTS method outperforms the two-stage FSM baseline while remaining worse than two of our proposed systems, ZMM-TTS2c and ZMM-TTS2x. We observed that the YourTTS model exhibits instability in its stochastic duration predictor, resulting in the production of unnatural durations for certain speakers and sentences.
On speaker similarity, ZMM-TTS2c and ZMM-TTS2x has better performance than YourTTS systems for the seen speaker condition in five languages (en, fr, ge, pt, and sw). However, under unseen conditions, YourTTS is better than our method in terms of DMOS in English, French, German and Spanish. This will be analyzed in detail in the next subsection.
@@ V-A4 Comparison between different input representations and languages
Interestingly, determining which is better for speech naturalness, character-based inputs or pre-trained phoneme representation based inputs, depends on the specific language and synthesis systems. FSM1x and FSM2x always have a higher MOS than FSM1c and FSM2c among the FSM-based systems. Utilizing pre-trained phoneme representations as input and XPhoneBERT as the text encoder can enhance the quality of synthesized speech. However, for our proposed ZMM-TTS system, better naturalness is not always achieved with ZMM-TTS1x and ZMM-TTS2x. For example, in Table VI, comparing the MOS of ZMM-TTS2x and ZMM-TTS2c, ZMM-TTS2x has achieved better results in English, French, and Portuguese, while ZMM-TTS2c has achieved better results in other languages. Furthermore, we observed that the MOS of ZMM-TTS2x is significantly better than ZMM-TTS2c in English. This may be because XPhoneBERT was pre-trained on a dataset with more English text than other languages. We will investigate how the amount of pre-training text affects XPhoneBERT’s representation performance in the future.
@@ V-A5 Comparison between seen and unseen speakers
An interesting result is that the naturalness of the synthesized speech of unseen speakers is not worse than that of seen speakers. However, there was still a gap between seen and unseen speakers in terms of DMOS for most languages. Although the pre-training process for the speaker encoder involves more than 7,000 speakers, the TTS system is only trained on a few hundred speakers. It is still challenging to achieve the ideal zero-shot performance when training with limited multilingual data.
@@ V-B Objective Evaluation Results
TABLE VII: CER (%) of ground-truth recordings (GT) and synthesized audio samples from various models.
@@ V-B1 SECS analysis
First, we found very strong correlations ($r=0.807$), using the Pearson correlation coefficient (PCC), between the SECS and the DMOS values. Compared with the FSM-based model, ZMM-TTS performs better in both SECS and DMOS metrics for both seen and unseen speakers. Compared with the YourTTS model, ZMM-TTS achieves similar SECS for seen speakers, but there are differences for unseen speakers, particularly English speakers. One possible reason for this difference is that the original implementation of YourTTS utilizes SCL to enable generalization in the characteristics of unseen speakers. The different SECS results of YourTTSE and YourTTSW also demonstrate the impact of SCL.
To investigate whether the difference is related to language and the impact of two speaker encoders, we plotted speaker embeddings of seen and unseen speakers from natural speech via T-SNE[82]. Figure 5 presents the T-SNE visualization and shows the following: Several unseen speakers are not close to the seen speakers in the training set. For example, there are two obvious outliers in the ECAPA-TDNN speaker embeddings of the English unseen speakers. We conducted a study to calculate the SECS values of the six unseen English speakers whose speech was synthesized by the ZMM-TTS1c system. We found two English outlier points in Figure 5(a) with SECS values of 0.766 and 0.644, which were significantly lower than that for the speaker who was closer to the seen English speakers. This explains why the speaker similarity of our proposed ZMM-TTS1c for English unseen speakers is significantly lower than that of YourTTS.
Although there are many different types of SOTA neural speaker embeddings, currently, ZMM-TTS has only attempted using a popular speaker representation from ECAPA-TDNN, and we will explore the performance with different speaker embeddings such as H/ASP and SCL in future work.
@@ V-B2 CER analysis
Table VII summarizes the obtained CER. The first finding is that the FSM1 and FSM2 baseline models perform better, although their sound quality is worse than the other systems. This demonstrates that the speech generated using Mel spectrograms as intermediate features is intelligible but not as natural sounding as when using SSL representations. Additionally, the speech synthesized through the Mel spectrograms tends to over-smooth the high-frequency region more than through SSL representations.
We also find that the CER is not always correlated with the MOS values. Although YourTTS CER is worse, the sound quality is obviously better compared with the FSM-based model. Furthermore, YourTTS CER in Portuguese and Swedish is significantly worse than in other languages. This also explains why the YourTTS system is significantly worse than our ZMM-TTS on MOS in these two languages. YourTTS only uses the character transcriptions rather than phonemes, which makes it more prone to mispronunciation issues as mentioned in [8].
Additionally, we found that the pre-trained phoneme representation input system performance is the best, and IPA is only better than characters in English. A notable result is that compared with FSM1, the CER performance of FSM2 has dropped significantly. For example, sometimes the pronunciation of the word “advice” in a sentence becomes incorrect and sounds more like the word “addressed,” resulting in ASR recognition errors. This shows that learning a discrete representation from Mel spectrograms to speech conversion in vec2wavVQ (Mel) can improve the quality of speech naturalness, but some fine-grained information related to linguistic content may be lost.
@@ VI Low-resource scenarios
TABLE VIII: Low-resource results. Total training data is 35 and 21 hours for Italian and Polish, respectively.
Italian Polish
Size Method UTMOS SECS CER MOS IMOS DMOS UTMOS SECS CER MOS IMOS DMOS
Few-shot scenarios
015m FSM2x 2.58 0.85 3.27 2.51±0.14 3.84±0.12 2.62±0.27 2.11 0.89 5.24 2.03±0.17 3.26±0.17 2.29±0.28
ZMM-TTS2x 2.97 0.91 3.92 3.33±0.17 4.09±0.13 2.95±0.28 2.77 0.93 8.27 3.54±0.23 3.49±0.18 2.91±0.30
FSM2c 2.10 0.78 7.13 1.62±0.14 3.02±0.13 2.29±0.25 1.45 0.83 28.36 1.41±0.16 1.53±0.17 1.81±0.23
ZMM-TTS2c 2.86 0.90 6.07 2.74±0.18 3.25±0.14 2.78±0.29 2.55 0.92 13.34 2.83±0.24 2.34±0.17 2.40±0.30
005m FSM2x 2.60 0.86 3.39 2.73±0.16 3.80±0.12 2.90±0.27 2.34 0.90 6.12 2.52±0.21 3.31±0.18 2.46±0.30
ZMM-TTS2x 2.92 0.90 4.02 3.09±0.17 3.88±0.13 2.78±0.29 2.69 0.92 10.01 3.28±0.23 3.28±0.20 2.84±0.32
2.5m FSM2x 2.48 0.84 3.12 2.44±0.14 3.62±0.12 2.75±0.26 2.22 0.89 11.29 2.05±0.20 2.56±0.18 2.28±0.30
ZMM-TTS2x 2.76 0.89 4.38 2.90±0.17 3.80±0.13 2.67±0.29 2.61 0.91 14.64 3.05±0.23 2.83±0.17 2.52±0.32
Zero-shot scenarios
$0$ FSM2x 2.27 0.73 4.10 1.66±0.15 3.30±0.14 1.43±0.15 2.33 0.77 15.93 1.60±0.18 1.80±0.17 1.34±0.16
ZMM-TTS2x 3.27 0.83 5.11 2.88±0.20 3.47±0.15 1.52±0.25 2.42 0.85 15.40 2.95±0.26 2.57±0.19 2.05±0.29
High-resource baseline
0001h YourTTS 1.64 0.74 89.77 - - - 1.40 0.65 88.86 - - -
35/21h YourTTS 2.91 0.96 6.01 - - - 1.70 0.71 89.68 - - -
35/21h FSM2x 2.27 0.91 3.29 2.25±0.12 4.21±0.14 3.53±0.26 2.00 0.88 2.89 2.08±0.17 3.79±0.21 2.58±0.31
35/21h ZMM-TTS2x 3.02 0.96 3.76 4.12±0.17 4.32±0.20 4.21±0.24 2.87 0.95 4.49 3.99±0.20 4.28±0.20 3.90±0.26
35/21h FSM2c 2.17 0.88 2.47 2.01±0.15 3.79±0.14 2.99±0.27 1.79 0.85 3.71 1.78±0.17 3.78±0.15 2.49±0.31
35/21h ZMM-TTS2c 2.92 0.96 2.22 3.77±0.17 4.29±0.12 4.19±0.24 2.70 0.95 2.42 4.04±0.17 4.29±0.14 3.93±0.26
- GT 2.94 0.99 3.17 4.58±0.14 4.80±0.08 4.32±0.24 2.82 0.99 3.95 4.64±0.12 4.88±0.08 4.50±0.18
In addition to the high-resource languages, we also evaluate the performance of the proposed method in unseen low-resource language TTS scenarios to investigate language adaptability with limited training data.
@@ VI-A2 Implementation Details
Considering the performance on six high-resource languages, we choose three models FSM2, ZMM-TTS2 and YourTTS to implement on new languages. We have attempted to use both characters and pre-trained phonemes as input representations. Low-resource language synthesis is investigated in two scenarios: few-shot and zero-shot. Few-shot means that the model is fine-tuned using the data from of the language before synthesing speech of that language. In contrast, as mentioned in Section I, zero-shot means that our model performs inference on unseen languages without model fine-tuning.
For the few-shot scenario, we created training sets of different sizes (2.5, 5, and 15 mins of audio). We use limited training data to finetune the model trained on six languages in Section IV. For language embeddings, we reserve a free ID for new languages, which will be trained from scratch using fine-tuning data only. Similarly, since different languages may use different characters, symbol embeddings for any characters unseen during pre-training must also be learned from fine-tuning data only. Specifically, in our experiments, we found that Italian and Polish had 3 and 14 unseen symbols, respectively, compared with the six pre-training languages. All few-shot models are fine-tuned for 10,000 steps with a batch size of 16. To compare the performance differences of the same language with high and low resources, we also perform fine-tuning with much larger datasets of 35h for Italian and 21h for Polish.
For the zero-shot scenario, to address the mismatch between language embeddings of seen and unseen languages, we removed the language embedding layer. Therefore, the models tested on the zero-shot scenario are trained as in Section IV on six seen languages but without a language embedding layer. Additionally, we used pre-trained phoneme representations as input, making it easily applicable to many unseen languages.
@@ VI-A3 Evaluation Methods
Objective Evaluation. We kept 100 sentences for the speech naturalness and speaker similarity tests and 1,000 sentences for ASR tests like those mentioned in Section IV-D. Furthermore, we adopted a publicly available automatic MOS (UTMOS) prediction model [84] to assess naturalness as in [5, 32].
Subjective Evaluation. We synthesized eight sentences from test sets for each TTS system. 15 native listeners of each language participated in our listening tests. First, we employed the same evaluation as described in Section IV-D to assess MOS and DMOS. Furthermore, ensuring the intelligibility of generated speech is crucial, particularly in low-resource scenarios. To this aim, we also designed an intelligibility test (IMOS) in which listeners provided ratings for speech regarding the extent to which words were pronounced wrong or were not understandable. The scale of IMOS is also from 1 (indicating very bad) to 5 (indicating very good).